OpenStudy (anonymous):

What are the characteristics of a radical equation? How is solving radical equations similar to solving linear equations? Why is it important to check the solutions to a radical equation? Create your own radical equation. Describe in complete sentences and demonstrate the process in finding its solution(s).

OpenStudy (lgbasallote):

uhhh 1) radical equations have...radicals -_- 2) you can substitute a radicand to x and solve like linear equations 3) because some solutions are imaginary (when it makes the radicand negative) 4) i suppose you can do that

OpenStudy (anonymous):

thanks youuuu

OpenStudy (lgbasallote):

by the way...i believe you are new to the site? it is usually customary here that askers award people with medals as a sign of appreciation and that they got what the answerer meant ;D but use it wisely and give it only to the person who gave you the answer that satisfied you the most

OpenStudy (anonymous):

\[\sqrt{8x+1=5}\] \[\sqrt{x-7+5=11}\] \[-4 \sqrt{x+9=20} \] \[\sqrt{6x+5=\sqrt{53}}\] \[\sqrt{x+2+4=x}\]

OpenStudy (anonymous):

I need solve, identify extraneous solutions, and show work

OpenStudy (lgbasallote):

the radical isnt very clear :/

OpenStudy (lgbasallote):

i mean the numbers that are included in the radical...it's not clear

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

Ok, i'm stupid at math... can you dumb that down for me..

OpenStudy (lgbasallote):

look at the picture or book you copied that from...do those look the same?

OpenStudy (anonymous):

Yepp, exactly the same.

OpenStudy (lgbasallote):

seriously?? there's a radical over the equal sign? o.O

OpenStudy (anonymous):

no, that is just over the equal sign cause of how the equation button is . so the radical is not over any of the equal signs.

OpenStudy (lgbasallote):

does it look like \[\sqrt{8x+1} = \sqrt 5\] or \[\sqrt{8x+ 1} = 5\]

OpenStudy (anonymous):

the second one